OpenStudy (anonymous):

A charge q of 7.3 × 1015 coulomb moves from point A to point B in an electric field. If the potential energy of the charge at point A is 3.5 × 10-12 joules and that at point B is 1.3 × 10-12 joules, what is the potential difference between point A and B?

OpenStudy (anonymous):

1.3 × 102 volts 3.0 × 102 volts 3.5 × 103 volts 6.8 × 103 volts

OpenStudy (farcher):

The charge is \[7.3 \times 10^{-15}\] coulombs not \[7.3 \times 10^{+15}\] coulombs. Potential difference is work done per unit charge and is measured in J/C = V. So you know the change in potential energy is the work done to move the charge and you know the charge so you can find the potential difference.